OpenStudy (askme12345):

Suppose your SSN is 121-21-2121. How many other people could have the same SSN? SSN(x)= (x-a1)(x+a2)(x-a3)(x+a4) and so on..

OpenStudy (anonymous):

What is a1, a2, a3,.... ?

OpenStudy (askme12345):

they are the numbers of the SSN for example if the number is 121-21-2121 in the SSN(x)- (x-1)(x+2)(x-1)(x+2) and so on

OpenStudy (anonymous):

Okay so you want a SSN such that SSN(x) = SSN(x') Now your SSN is 121212121 4 2s and 5 1s and In total 9 numbers

OpenStudy (anonymous):

Do you know Permutations?

OpenStudy (askme12345):

barely sorry

OpenStudy (anonymous):

Hmm I am not sure, Maybe \[\Huge9_{P_{4}} + 9_{P_{5}}\] \[\frac{9!}{5!} + \frac{9!}{4!}\] \[ 9\times8\times7\times6 + 9\times8\times7\times6\times5\] People can have same SSN polynomial as yours

OpenStudy (askme12345):

i have to state one that would be? how can i find one that is equ to 121212121

OpenStudy (anonymous):

Your SSN is unique but not your SSN polynomial

OpenStudy (askme12345):

yeah I understand I have to find an SSN polynomial equal to that ssn

OpenStudy (turingtest):

Oh no! Not this problem again. Askme, Your teacher is a sadist!

OpenStudy (anonymous):

wtf is ssn

OpenStudy (turingtest):

Social Security Number lol!